OpenStudy (anonymous):

write the linear function where f(-3)=14 and f(5)=-10

OpenStudy (mathmate):

It means find the equation of a line (linear function) that passes through (-3,14) and (5,-10).

OpenStudy (anonymous):

haha i know that i dont know what the equation is

OpenStudy (anonymous):

y – y1 = m(x – x1) m = slope

OpenStudy (anonymous):

1. FInd the slope, write it down. 2. choose one pair of coordinates and plug them into y – y1 = m(x – x1) y - 14 = m(x - -3) or y - - 10 = m(x - 5)

OpenStudy (anonymous):